# Spatially Resolved Kinematics of Post-starburst Galaxies

### v1

@@ 4.2 Evolution of angular momentum with asymmetry
Figure 6: Angular momentum at half-light radius, $\lambda_{Re}$, versus asymmetry, $A$, for 92 PSBs (black points), 1352 SFs (blue contours), and 903 ETGs (red contours) from the MaNGA survey. The black line represents the running average for the PSBs. The ETGs tend to be fast and slow rotators for similar asymmetry values, while angular momentum is correlated with asymmetry for the SFs. The PSBs lie between the ETGs and SFs in both angular momentum and asymmetry.We do not observe a significant trend between angular momentum and asymmetry within the PSB sample. Thus, if PSBs are more likely to be slow rotators after a major merger, the rapid decay of observable tidal features may have obscured this effect.
@@
Asymmetric features in galaxies can be used to trace their recent merger history due to the presence of tidal tails, shells, and companion galaxy remnants, as well as to separate galaxies between morphological types. The rotational asymmetry $A$, a purely morphological parameter, is calculated by rotating a galaxy by 180o and subtracting it from the original image, and further applying relevant corrections to take into account background noise etc.
@@
PSBs tend to exhibit high asymmetry compared to the ETGs, which could be due to tidal/relic features associated with the recent starburst and/or recent interaction with another galaxy or dust lanes in the case of dusty PSBs (Yang et al., 2008; Sazonova et al., 2021). The SFs have high $A$ due to the presence of spiral arms and flocculence in the disk. However, the ETGs tend to have low $A$ due to the absence of tidal features or spiral arms.
@@
The asymmetry, thus, provides a way to test for the effects of recent mergers and interaction in galaxies and test for trends with angular momentum content. We would expect the PSBs resulting from major mergers to be characterized by significantly lower angular momentum, however, the expected trends depend on the relative timescales for tidal features to be present and angular momentum to be lost. Low values of $A$ could be present in post-major-merger galaxies due to the fact that their merger features fade quickly as the PSBs age (Pawlik et al., 2016). The PSBs also have a large range of angular momentum, which can obscure any subtle trends we might observe.
@@
Asymmetry values for the PSBs, ETGs, and SFs in the MaNGA survey are obtained from the MaNGA Visual Morphology Catalogue (Vázquez-Mata et al., 2022). Following digital image post-processing, the catalogue detects the inner structures and external, low surface brightness characteristics. The morphological classification process was implemented empirically, based on the methods described in Hernández-Toledo et al. (2010) and Cheng et al. (2011). They reported an estimation of the CAS parameters (Concentration, Asymmetry, and Clumpiness) derived from the Legacy Surveys and SDSS imaging, following the method by Conselice (2003).
@@
Figure 6 compares the angular momentum against rotational asymmetry for 92 PSBs, 1357 SFs, and 903 ETGs from the MaNGA survey. The black line is the running average for the PSBs. We observe that the slow and fast rotators in the ETG sample are concentrated around small values of asymmetry, however, the fast rotators in the SF sample tend to exhibit high asymmetry values. The finding of higher angular momentum in the SF galaxies with more asymmetry may be due to a similar effect as seen by Romanowsky & Fall (2012). They find that later Hubble-type galaxies and galaxies with lower bulge-to-total light ratios have more angular momentum, even within the class of spirals. The PSBs, on the other hand, do not seem to exhibit significant correlation (Spearman correlation: $\rho$ = 0.19, p = 0.11) for angular momentum and asymmetry. Thus, if PSBs are more likely to be slow rotators after a major merger, the rapid decay of observable tidal features may have obscured this effect.
@@
Deeper photometry may reveal more subtle trends between the asymmetry caused by merger features and the effect on angular momentum. Rutherford et al. (2024) used deep imaging from Subaru-Hyper Suprime Cam to test for the presence of tidal features and shells around galaxies as a function of angular momentum, finding a correlation between the presence of shells and lower angular momentum in a subset of the younger galaxies. Deeper photometry in next-generation surveys such as the Rubin Observatory LSST will provide useful constraints on the correlation of merger features with angular momentum during the PSB phase and into the ETG phase.
@@ 5 Conclusions
We trace the evolution of a galaxy as it evolves from the post-starburst phase to the quiescent phase by comparing and contrasting spatially resolved properties of 92 post-starbursts from the MaNGA survey with 260 early-type galaxies from the ATLAS3D project, 1577 MaNGA star-forming galaxies, and 964 early-types from MaNGA. We attempt to glean information about the merger history of these galaxies by studying their inner kinematics and classifying them as slow and fast rotators based on their angular momentum and ellipticity at the effective radius. Our conclusions are as follows:
@@
    We find that the PSB sample has fewer slow rotators than the ETG samples from both ATLAS3D and the MaNGA DR14 (Graham et al., 2018), with the PSB sample consisting of $\approx 5\%$ (5/92), the ATLAS3D ETG sample having $\approx 14\%$ (36/260), SFs from MaNGA having $\approx 4\%$ (64/1577), and MaNGA ETGs having $\approx 22\%$ (213/964) slow rotators. Star-forming galaxies from Graham et al. (2018) have consistently high values of angular momentum, with a mean value above even the fastest rotator PSBs in our sample.
@@
    We note that the angular momentum is generally lower for ETGs with higher stellar mass, which is not the case for the PSBs or SFs. This is evident from our finding that $\sim 58\%$ of slow rotator MaNGA ETGs and $58\%$ of slow rotator ATLAS3D ETGs are in the top 30th percentile of the respective ETG stellar mass distributions, while only $40\%$ slow-rotating MaNGA PSBs and $42\%$ slow-rotating MaNGA SFs are in the top 30th percentile of the respective PSB and SF stellar mass distribution.
@@
    From our analysis of the angular momentum in the context of mass-weighted and light-weighted ages, we find that the angular momentum of SFs and ETGs is more tightly correlated with MW age as compared to the PSBs. We observe a significant decrease in angular momentum with MW age for both the ETG and SF samples, though not within the PSB sample. For LW age, we only observe a significant correlation within the ETG sample.
@@
    Our analysis of galaxy kinematics with respect to asymmetry does not suggest a clear trend for the PSBs (Spearman correlation: $\rho$ = 0.19, p = 0.11). Both slow and fast rotator ETGs have similarly low values of asymmetry, while the SF sample displays a correlation of higher asymmetry in the highest angular momentum galaxies.
This work constrains the possible formation and evolution of ETGs. Roughly half of ETGs are expected to have gone through a PSB phase after rapidly quenching (Wild et al., 2020), so if slow rotator ETGs are created through gas-rich merger mergers, we should observe a similar or greater fracion of slow rotators in the PSBs. However, we find more slow rotators in the ETG sample as compared to the PSBs. The tight correlation of angular momentum with the mass-weighted age, as compared to the luminosity-weighted age, indicates that angular momentum is more closely coupled to the longer-scale formation history of the galaxy, instead of the more recent history traced by the light-weighted age. PSBs already have lower angular momentum than SF/disk galaxies, and since kinematic evolution takes place over the many-Gyr lifetime of a galaxy, PSBs must still lose significant angular momentum before evolving to ETGs. This is consistent with the trends in angular momentum with stellar mass, which imply that the slow rotator ETGs have gained mass and lost angular momentum over many Gyr since they last formed stars in situ or experienced a starburst. The long timescale of kinematic evolution is probably why we do not observe significant trends for angular momentum with asymmetry as the asymmetry varies quite rapidly during/after a merger. These implications suggest that the PSBs will undergo multiple major/minor dry merger events as they evolve to the ETG phase while shedding off their angular momentum. This scenario is consistent with the findings of simulations presented in Naab et al. (2014), where the slow rotator ETGs tended to have many dry mergers over being quiescent for a long time.
@@ Acknowledgements
We thank Akshat Tripathi for useful discussions. This work was made possible through the Preparing for Astrophysics with LSST Program, supported by the Heising-Simons Foundation and managed by Las Cumbres Observatory.
@@ Appendix A Comparison of Photometry
In order to analyze the kinematics within 1 Re, we need to know the shape and size of the galaxy. We also need the Sérsic index to estimate the impact of beam smearing on the photometry. We compare the photometric properties from the NSA catalog with the ones given in Table 3 of Simard et al. (2011). Simard et al. (2011) performed bulge+disk decompositions in g- and r-band by convolving the point spread function in 2D for 1,123,718 galaxies from the Legacy portion of SDSS Release 7. They use three fitting models - a pure Sérsic model, an nb = 4 bulge + disk model, and a Sérsic (free nb ) bulge + disk model. We chose the data calculated using the pure Sérsic model as NSA also used a Sérsic fit, making the two sets of photometric properties comparable. We compare the galaxy Sérsic index (ng), the galaxy position angle (phi), galaxy ellipticity (e), and the galaxy half-light radius in r-band (Rhlr).
@@
To test the consistency of the photometry across all the galaxies, and not just the PSBs, we plot the photometric properties from the NSA catalog against the respective properties from Simard et al. (2011), as shown in Figure 7. The gray contours represent all the galaxies in both catalogs, and the blue contour lines represent the 92 PSBs across the two catalogs. We observe that the scatter in the photometric properties of the PSBs around the red 1:1 line is quite consistent with the overall scatter of all the galaxies for all the properties. This means that overall, we do not expect a particular bias for the PSBs, although the scatter in properties can be large.
@@
To further test our choice to use the NSA photometry, we compare the NSA catalog to the MaNGA PyMorph DR17 photometric catalogue. PyMorph provides photometric properties determined by fitting Sérsic and Sérsic+Exponential models to the 2D surface brightness profiles of galaxies in the MaNGA DR17 release (Meert et al., 2013; Fischer et al., 2017; Bernardi et al., 2017; Domínguez Sánchez et al., 2021). PyMorph was re-run to incorporate the SDSS DR14 images, improved bulge-to-disk decomposition, and all the fits were re-fit if deemed necessary (Fischer et al., 2018; Domínguez Sánchez et al., 2021). Since NSA performed a pure Sérsic fit, we selected the PyMorph photometry obtained from fitting the Sérsic model. We compare the Sérsic index (N_S), photometric angle (90 - PA_S - SPA), ellipticity (1-BA_S), and effective radius (obtained from the semi-major axis, A_hl_S) in r-band from PyMorph runs with the corresponding properties in r-band from the NSA catalog, as shown in Figure 8. The gray contours represent all the galaxies in both catalogs, and the blue contour lines represent the 92 PSBs across the two catalogs. Unlike the case for the Simard et al. (2011) parameters, the scatter in the PSBs is larger than the typical scatter for normal galaxies, which indicates the parameters derived for the PSBs may contain significant outliers.
@@
Figure 7: The four panels compare the Sérsic index (top left), photometric angle (top right), ellipticity (bottom left), and half-light radius (bottom right) of all the galaxies from the NSA catalog with the properties given in Table 3 of Simard et al. (2011), all of which were determined using a pure Sérsic model. The gray shaded contour region corresponds to all the galaxies while the blue contour lines represent the 92 PSBs. We use the red 1:1 line to check the consistency of the scatter across the PSB sample with all the galaxies in the two catalogs. We observe that the scatter in the photometric properties of the PSBs around the red 1:1 line is quite consistent with the overall scatter of all the galaxies for all the properties. This means that overall, we do not expect a particular bias for the PSBs, although the scatter in properties can be large.
The scatter observed in photometry properties between these catalogs may mean that the choice of catalog influences our measurement of specific angular momentum. The $\lambda_{Re}$ obtained using the three catalogs are plotted against each other in Figure 9. We do not observe any particular trend in the scatter and instead find that the points tend to scatter around the 1:1 black line, proving that choosing another catalog would not affect our results and conclusions drastically. We find 5.4% (5/92) slow rotators using the NSA catalog, 7.6% (7/92) slow rotators using Simard et al. (2011) photometric properties, and 9.8% (9/92) slow rotators using the PyMorph catalogue. In all three cases, the PSBs have significantly fewer slow rotators than the ETG samples.
As a further test, we over-plot the SDSS images of the PSBs with ellipses created using the photometric properties of the respective catalogs. We observe that the ellipse corresponding to NSA photometry aligns quite well with the galaxy, however, the ellipses traced by photometric properties from Simard et al. (2011) and the PyMorph catalogue do not align as well, and in some cases are aligned away from the galaxy by $\sim 90^{o}$ for Simard et al. (2011) photometry. Thus, we use the NSA catalog to measure $\lambda_{Re}$ in this work as described in §3.1.
@@
Figure 8: The four panels compare the Sérsic index (top left), photometric angle (top right), ellipticity (bottom left), and half-light radius (bottom right) of all the galaxies from the NSA catalog with the corresponding pure Sérsic model properties provided in MaNGA PyMorph DR17 photometric catalogue. The gray shaded contour region corresponds to all the galaxies while the blue contour lines represent the 92 PSBs. We use the red 1:1 line to check the consistency of the scatter across the PSB sample with all the galaxies in the two catalogs. Unlike the case for the Simard et al. (2011) parameters, the scatter in the PSBs is larger than the typical scatter for normal galaxies, which indicates the parameters derived for the PSBs may contain significant outliers.
Figure 9: Comparison of the angular momentum obtained using photometry from the NSA catalog, Simard et al. (2011) and MaNGA PyMorph DR17 photometric catalogue for the PSBs. We do not observe a systematic scatter as the points tend to scatter about the 1:1 black line. We find 5.4% (5/92) slow rotators using the NSA catalog, 7.6% (7/92) slow rotators using Simard et al. (2011) photometric properties, and 9.8% (9/92) slow rotators using the PyMorph catalogue. In all three cases, the PSBs have significantly fewer slow rotators than the ETG samples.

### v2

@@ 4.2 Evolution of Angular Momentum with Asymmetry
Figure 6: Angular momentum at half-light radius, $\lambda_{Re}$, vs. asymmetry, $A$, for 79 PSBs (black points), 1341 SFGs (blue contours), and 841 ETGs (red contours) from the MaNGA survey. The black line represents the running average for the PSBs. The ETGs tend to be fast and slow rotators for similar asymmetry values, while angular momentum is correlated with asymmetry for the SFGs. The PSBs lie between the ETGs and SFGs in both angular momentum and asymmetry. We do not observe a significant trend between angular momentum and asymmetry within the PSB sample. Thus, if PSBs are more likely to be slow rotators after a major merger, the rapid decay of observable tidal features may have obscured this effect.
@@
Asymmetric features in galaxies can be used to trace their recent merger history due to the presence of tidal tails, shells, and companion galaxy remnants, as well as to separate galaxies between morphological types. The rotational asymmetry $A$, a purely morphological parameter, is calculated by rotating a galaxy by 180 and subtracting it from the original image, and further applying relevant corrections to take into account background noise, etc.
@@
PSBs tend to exhibit high asymmetry compared to the ETGs, which could be due to tidal/relic features associated with the recent starburst and/or recent interaction with another galaxy or dust lanes in the case of dusty PSBs (Yang et al., 2008; Sazonova et al., 2021). The SFGs have high $A$ due to the presence of spiral arms and flocculence in the disk. However, the ETGs tend to have low $A$ due to the absence of tidal features or spiral arms.
@@
The asymmetry thus provides a way to test for the effects of recent mergers and interaction in galaxies, and test for trends with angular momentum content. We would expect the PSBs resulting from major mergers to be characterized by significantly lower angular momentum, however the expected trends depend on the relative timescales for tidal features to be present and angular momentum to be lost. Low values of $A$ could be present in post-major-merger galaxies due to the fact that their merger features fade quickly as the PSBs age (Pawlik et al., 2016). The PSBs also have a large range of angular momentum, which can obscure any subtle trends we might observe.
@@
Asymmetry values for the PSBs, ETGs, and SFGs in the MaNGA survey are obtained from the MaNGA Visual Morphology Catalogue (Vázquez-Mata et al., 2022). Following digital image postprocessing, the catalogue detects the inner structures and external, low-surface-brightness characteristics. The morphological classification process was implemented empirically, based on the methods described in Hernández-Toledo et al. (2010) and Cheng et al. (2011). They reported an estimation of the CAS parameters (concentration, csymmetry, and clumpiness) derived from the Legacy Surveys and SDSS imaging, following the method by Conselice (2003).
@@
Figure 6 compares the angular momentum against rotational asymmetry for 78 PSBs, 1341 SFGs, and 841 ETGs from the MaNGA survey. The black line is the running average for the PSBs. We observe that the slow and fast rotators in the ETG sample are concentrated around small values of asymmetry, however the fast rotators in the SFG sample tend to exhibit high asymmetry values. The finding of higher angular momentum in the SFGs with more asymmetry may be due to a similar effect as seen by Romanowsky & Fall (2012). They find that later Hubble-type galaxies and galaxies with lower bulge-to-total light ratios have more angular momentum, even within the class of spirals. The PSBs, on the other hand, do not seem to exhibit a correlation (Spearman correlation: $\rho\approx$ 0.15, p $\approx$ 0.21) for angular momentum and asymmetry. Thus, if PSBs are more likely to be slow rotators after a major merger, the rapid decay of observable tidal features may have obscured this effect.
@@
Deeper photometry may reveal more subtle trends between the asymmetry caused by merger features and the effect on angular momentum. Rutherford et al. (2024) used deep imaging from Subaru Hyper Suprime-Cam to test for the presence of tidal features and shells around galaxies as a function of angular momentum, finding a correlation between the presence of shells and lower angular momentum in a subset of the younger galaxies. Deeper photometry in next-generation surveys such as the Rubin Observatory LSST will provide useful constraints on the correlation of merger features with angular momentum during the PSB phase and into the ETG phase.
4.3 Testing for Selection Bias in the PSB Sample
We consider the range in properties of the MaNGA PSBs compared to the parent PSB sample in order to assess biases in this sample that may affect our conclusions. The MaNGA survey targets low-redshift galaxies, so we must ensure this does not result in biases in the trends with angular momentum. In Figure 7, we compare the redshift and stellar mass of the 89 MaNGA PSBs (black points) to those of the parent sample of $\sim$5000 PSBs (green contours) from French et al. (2023). While the redshift cut of MaNGA is apparent, and the MaNGA PSBs have systematically lower stellar masses as a result, we see that the MaNGA PSBs nonetheless span a wide range in stellar mass. Furthermore, when we consider the trends in stellar mass and angular momentum in Figure 4, we see that the PSBs have higher $\lambda_{Re}$ than the ETG sample over a wide range in stellar mass. Thus, the lack of higher stellar mass PSBs at larger distance does not hinder our primary conclusions. When we compare the range of other quantities, namely the stellar age as traced by Dn4000 and ellipticity, a Kolmogorov-Smirnov test shows that we cannot exclude the null hypothesis that the MaNGA PSBs are consistent with the parent sample of SDSS PSBs.
Figure 7: We test for dependence between redshift and stellar mass by comparing the smaller PSB sample used in this work with the combined PSB sample from French et al. (2023). The black points are the 89 MaNGA PSBs while the green contours represent the 4985 PSBs from the combined sample. We find that despite the PSBs preferring lower redshifts, they span the entire stellar mass range of the bigger sample.
@@ 5 Conclusions
We trace the evolution of a galaxy as it evolves from the post-starburst phase to the quiescent phase by comparing and contrasting spatially resolved properties of 89 post-starbursts (PSBs) from the MaNGA survey with 260 early-type galaxies (ETGs) from the ATLAS3D project, 1592 MaNGA star-forming galaxies (SFGs), and 916 ETGs from MaNGA. We attempt to glean information about the merger history of these galaxies by studying their inner kinematics and classifying them as slow and fast rotators based on their angular momentum and ellipticity at the effective radius. Our conclusions are as follows:
@@
    We find that the PSB sample has fewer slow rotators than the ETG samples from both ATLAS3D and MaNGA, with the PSB sample consisting of $\sim 6\%$ (5/89), the ATLAS3D ETG sample having $\sim 14\%$ (36/260), and MaNGA ETGs having $\sim 19.5\%$ (196/916) slow rotators. The SFGs from MaNGA have fewer slow rotators ($\sim 3.5\%$, 56/1592) than the PSB and ETG samples, with the median angular momentum of the SFG sample higher than that of the fast-rotator PSB and ETG samples.
@@
    We note that the angular momentum is generally lower for ETGs with higher stellar mass, which is not the case for the PSBs or SFGs. This is evident from our finding that $\sim 58\%$ of slow-rotator MaNGA ETGs and $\sim 58\%$ of slow-rotator ATLAS3D ETGs are in the top 30th percentile of the respective ETG stellar mass distributions, while only $40\%$ slow-rotating MaNGA PSBs and $\sim 32\%$ slow-rotating MaNGA SFGs are in the top 30th percentile of the respective PSB and SFG stellar mass distribution.
@@
    From our analysis of the angular momentum in the context of mass-weighted (MW) and luminosity-weighted (LW) ages, we find that the angular momentum of SFGs and ETGs is more tightly correlated with MW and LW ages as compared to the PSBs. We observe a significant decrease in angular momentum with age for the ETG and SFG samples, though not within the PSB sample.
@@
    Our analysis of galaxy kinematics with respect to asymmetry does not suggest a clear trend for the PSBs (Spearman correlation: $\rho\approx$ 0.15, p $\approx$ 0.21). Both slow- and fast-rotator ETGs have similarly low values of asymmetry, while the SFG sample displays a correlation of higher asymmetry in the highest-angular-momentum galaxies.
This work constrains the possible formation and evolution of ETGs. Roughly half of the ETGs are expected to have gone through a PSB phase after rapidly quenching (Wild et al., 2020), so if slow-rotator ETGs are created through gas-rich mergers, we should observe a similar or greater fraction of slow rotators in the PSBs. However, we find more slow rotators in the ETG sample as compared to the PSBs. The tighter anticorrelation of angular momentum with the MW and LW age for the SFGs and ETGs, as compared to the PSBs, indicates that angular momentum is more closely coupled to the longer-scale formation history and evolution of the galaxy while it is in the SFG and ETG phases, instead of the time it spends in the green valley region. PSBs already have lower angular momentum than star-forming/disk-like galaxies, and since kinematic evolution takes place over the multi-Gyr lifetime of a galaxy, PSBs must still lose significant angular momentum before evolving to ETGs. This is consistent with the trends in angular momentum with stellar mass, which imply that the slow-rotator ETGs have gained mass and lost angular momentum over many billion years since they last formed stars in situ or experienced a starburst. The long timescale of kinematic evolution is probably why we do not observe significant trends for angular momentum with asymmetry as the asymmetry varies quite rapidly during/after a merger. These implications suggest that the PSBs will undergo multiple major/minor dry merger events as they evolve to the ETG phase while shedding off their angular momentum. This scenario is consistent with the findings of simulations presented in Naab et al. (2014), where the slow-rotator ETGs tended to have many dry mergers while being quiescent for a long time.
@@ Acknowledgements
We thank the referee for their comments, which have improved this manuscript. We thank Akshat Tripathi for useful discussions. This work was made possible through the Preparing for Astrophysics with LSST Program, supported by the Heising-Simons Foundation and managed by Las Cumbres Observatory.
@@ Appendix A Comparison of Photometry
In order to analyze the kinematics within 1 $Re$, we need to know the shape and size of the galaxy. We also need the Sérsic index to estimate the impact of beam smearing on the photometry. We compare the photometric properties from the NSA catalog with the ones given in Table 3 of Simard et al. (2011). Simard et al. (2011) performed bulge+disk decompositions in the $g$ and $r$ bands by convolving the PSF in 2D for 1,123,718 galaxies from the Legacy portion of SDSS Data Release 7. They use three fitting models: a pure Sérsic model, an nb = 4 bulge + disk model, and a Sérsic (free nb) bulge + disk model. We chose the data calculated using the pure Sérsic model as NSA also used a Sérsic fit, making the two sets of photometric properties comparable. We compare the galaxy Sérsic index (ng), the galaxy position angle (phi), galaxy ellipticity (e), and the galaxy half-light radius in the $r$ band (Rhlr).
@@
To test the consistency of the photometry across all the galaxies, and not just the PSBs, we plot the photometric properties from the NSA catalog against the respective properties from Simard et al. (2011), as shown in Figure 8. The gray contours represent all the galaxies obtained by cross-matching MaNGA with Simard et al. (2011) data irrespective of galaxy type, and the blue points represent the 84 PSBs with photometric data available in Simard et al. (2011). We observe that the scatter in the photometric properties of the PSBs around the red 1:1 line is quite consistent with the overall scatter of all the galaxies for all the properties, though there are some systematic offsets in photometric angle and effective radius.
@@
To further test our choice to use the NSA photometry, we compare the NSA catalog to the MaNGA PyMorph DR17 photometric catalogue. PyMorph provides photometric properties determined by fitting Sérsic and Sérsic+Exponential models to the 2D surface brightness profiles of galaxies in the MaNGA DR17 release (Meert et al., 2013; Fischer et al., 2017; Bernardi et al., 2017; Domínguez Sánchez et al., 2021). PyMorph was rerun to incorporate the SDSS DR14 images, improved bulge-to-disk decomposition, and all the fits were re-fit if deemed necessary (Fischer et al., 2018; Domínguez Sánchez et al., 2021). Since NSA performed a pure Sérsic fit, we selected the PyMorph photometry obtained from fitting the Sérsic model. We compare the Sérsic index (N_S), photometric angle (90 - PA_S - SPA), ellipticity (1-BA_S), and effective radius (obtained from the semi-major axis, A_hl_S) in the $r$ band from PyMorph runs with the corresponding properties in the $r$ band from the NSA catalog, as shown in Figure 9. The gray contours represent all the galaxies obtained by cross-matching MaNGA with the PyMorph catalogue irrespective of phase, and the blue points represent the 89 PSBs across the two catalogs. Similar to the comparison between the MaNGA and Simard et al. (2011) data, we observe similarities in the PSB sample with the overall galaxy sample, as well as systematic offsets and scatter. The impact of this offset and scatter in the photometric properties on $\lambda_{Re}$ is further explored in Figure 10.
@@
Figure 8: The four panels compare the Sérsic index (top left), photometric angle (top right), ellipticity (bottom left), and half-light radius (bottom right) of all the galaxies from the NSA catalog with the properties given in Table 3 of Simard et al. (2011), all of which were determined using a pure Sérsic model. The gray shaded contour region corresponds to all MaNGA galaxies cross-matched with the Simard et al. (2011) data, while the blue points represent the 84 MaNGA PSBs. We use the red 1:1 line to check the consistency of the scatter across the PSB sample with all the galaxies in the two catalogs. We find that the PSBs are similarly distributed as the total galaxy sample, though we observe significant scatter, and some systematic offsets. In Figure 10, we test the impact of these offsets on our measurement of $\lambda_{Re}$.
@@
Figure 9: The four panels compare the Sérsic index (top left), photometric angle (top right), ellipticity (bottom left), and half-light radius (bottom right) of all the galaxies from the NSA catalog with the corresponding pure Sérsic model properties provided in MaNGA PyMorph DR17 photometric catalogue. The gray shaded contour region corresponds to all MaNGA galaxies cross-matched with the PyMorph catalogue, while the blue points represent the 89 PSBs. We use the red 1:1 line to check the consistency of the scatter across the PSB sample with all the galaxies in the two catalogs. We find that the PSBs are similarly distributed as the total galaxy sample, though we observe significant scatter, and some systematic offsets. In Figure 10, we test the impact of these offsets on our measurement of $\lambda_{Re}$.
Figure 10: We test whether our results are consistent with respect to photometry by comparing the results from the NSA photometric properties with that of MaNGA PyMorph DR17 photometric catalogue and Simard et al. (2011) photometric data. Left: comparison of the angular momentum obtained using photometry from the NSA catalog and Simard et al. (2011), in conjunction with MaNGA DR17 maps for 81 PSBs (black points), 1512 SFGs (blue contours) and 831 ETGs (red contours). The $\lambda_{Re}$ values determined using Simard et al. (2011) data tend to be lower in comparison to the predictions from NSA photometry, and there is significant scatter about the 1:1 black line, especially for higher values of $\lambda_{Re}$. Using data from Simard et al. (2011), we find $\sim 5\%$ slow rotators in the SFGs, $\sim 7.4\%$ slow rotators in the PSBs and $\sim 14.8\%$ slow rotators in the ETG sample. Right: we compare NSA with PyMorph photometry, while using MaNGA DR17 maps in our calculation of $\lambda_{Re}$ for 83 PSBs (black points), 1502 SFGs (blue contours) and 853 ETGs (red contours). We observe even less scatter about the 1:1 black line as compared to the left panel, and PyMorph predicts smaller values of $\lambda_{Re}$ as compared to NSA as well. Using PyMorph photometric properties, we find $\sim 5.7\%$ slow rotators in the SFGs, $\sim 8.4\%$ slow rotators in the PSBs and $\sim 19.1\%$ slow rotators in the ETG sample. Irrespective of the catalog being used, the PSBs have fewer slow rotators than the ETGs and tend to rotate faster than the SFGs. This confirms that the choice of photometry would not have a significant impact on our overall results.
The scatter observed in photometric properties between these catalogs may mean that the choice of catalog influences our measurement of specific angular momentum. The $\lambda_{Re}$ obtained using the three photometric catalogs, MaNGA DR17 data, and the method we describe above are plotted against each other in Figure 10. The Sérsic indices were modified as explained in Section 3.1 to ensure that they are in the range $0.5-6.5$ for 11/84 PSBs from the Simard et al. (2011) sample and 15/89 PSBs from the PyMorph catalogue. We cannot accurately measure $\lambda_{Re}$ for 3/84 PSBs as their semimajor axis from Simard et al. (2011) data is smaller than their $\sigma_{PSF}$, so we only include 81/84 PSBs in Figure 10(a). Similarly, Figure 10(b) consists of only 83/89 PSBs due to missing $\epsilon_{Re}$ measurements for the excluded galaxies in PyMorph photometry. The $\lambda_{Re}$ values predicted using PyMorph agree better with the predictions from NSA photometric data as compared to the Simard et al. (2011) data. We find that, on average, Simard et al. (2011) and PyMorph catalogs predict lower $\lambda_{Re}$ as compared to NSA. Using Simard et al. (2011) photometry, we find $\sim 5\%$ slow rotators in the SFGs, $\sim 7.4\%$ slow rotators in the PSBs, and $\sim 14.8\%$ slow rotators in the ETGs. Similarly, we find $\sim 5.7\%$ slow rotators in the SFGs, $\sim 8.4\%$ slow rotators in the PSBs, and $\sim 19.1\%$ slow rotators in the ETG sample using PyMorph photometry. NSA photometry predicts the least number of slow rotators for the SFGs and PSBs as compared to the Simard et al. (2011) and PyMorph catalogs. However, the PSBs have fewer slow rotators than the ETGs and rotate faster as compared to the SFGs irrespective of the photometry utilized, which confirms that choosing another photometric catalog would not affect our qualitative results and conclusions drastically. Thus, the NSA catalog is a suitable choice for measuring $\lambda_{Re}$ in this work as described in Section 3.1.
Figure 11: We check for the impact of the MaNGA data release used on our results by drawing comparisons between DR14 and DR17 data using photometry from Graham et al. (2018) and NSA. Left: we compare the $\lambda_{Re}$ obtained from photometric properties determined by Graham et al. (2018) and DR14 with the values obtained using DR17 in conjunction with NSA photometry. The black points are the 16 MaNGA PSBs, the blue contours are the 1571 SFGs and the red contours are the 914 ETGs that are present in both the photometric catalogs and MaNGA data releases. We find that the $\lambda_{Re}$ values are scattered around the 1:1 (black dashed) line, with the scatter increasing for higher values of $\lambda_{Re}$ for all three classes. Right: to analyze how the difference in the data release affects our results, we set the data release to DR17 and calculate $\lambda_{Re}$ using Graham et al. (2018) and NSA photometry respectively for 16 PSBs, 1569 SFGs and 914 ETGs present in both of the photometric catalogs. We observe a significant reduction in scatter around the 1:1 line for all three classes, owing to the most recent data release being utilized. As evident from the two panels, the MaNGA data release used for obtaining the kinematic spectra has a non-negligible impact on the kinematic analysis, which contains key implications for the overall kinematic evolution of the galaxy.
We perform a similar test on the different data releases of MaNGA to compare directly with the measurements from Graham et al. (2018) and check for any systematic differences. We compare MaNGA DR14 and DR17 data along with photometric properties from Graham et al. (2018) and the NSA catalog while uniformly using the function given by Graham et al. (2018) for calculation of $\lambda_{Re}$ in Figure 11. In Figure 11(a), we compare the $\lambda_{Re}$ from Graham et al. (2018) (which uses MaNGA DR14 and photometric properties derived in that work) and the $\lambda_{Re}$ we recalculate using DR17 and the NSA photometry. As evident from Figure 11(a), the SFGs, PSBs and ETGs seem to scatter around the 1:1 line (black dashed line). We also observe systematic shifts that differ between samples. The scatter in both the samples increases with $\lambda_{Re}$. In Figure 11(b), we compare the $\lambda_{Re}$ calculated using photometric properties derived by Graham et al. (2018) with NSA photometry while setting the data release to DR17 for both. Fixing the data release to DR17 and only varying the photometric properties results in a significant reduction in the scatter in $\lambda_{Re}$ for the SFGs, PSBs and ETGs. This shows that the data release being utilized in the calculation for $\lambda_{Re}$ has a significant impact on our results, more so than the photometric parameters used.
Because of the improvements in DR17 that we discuss in Section 2.2, we use our rereduced $\lambda_{Re}$ values (the y-axis of Figures 11(a) and (b)) for the SFG and ETG samples in this manuscript. The method for calculating $\lambda_{Re}$ is thus the same for both the PSB and comparison samples.